OpenStudy (anonymous):

Solve Each Equation by first clearing the fractions or decimals 1/2(4r+1) -r=2/5(2r-3)+2/3 I Need to show all work

OpenStudy (anonymous):

\[\frac{ 1 }{ 2 }(4r+1)-r=\frac{ 2 }{ 5 }(2r-3)+\frac{ 3 }{ 2 }\]

OpenStudy (mertsj):

Common denominator is?

OpenStudy (anonymous):

10

OpenStudy (mertsj):

Multiply each term by 10

OpenStudy (mertsj):

\[\frac{4r+1}{2}\times 10-r \times10=\frac{2(2r-3)}{5}\times 10=\frac{3}{2}\times10\]

OpenStudy (anonymous):

ok i got \[\frac{ 5r }{ 2 }\times 10-r \times 10= \frac{ 4r-3 }{ 5 } \times 10= \frac{ 3 }{ 2 } \times \frac{ 10 }{ 1 }\]